Pipe A can fill a tank in 8 hours. When pipe A and pipe B are used, the tank is filled in 5 hours. How long, in hours, would it take pipe B alone to fill the tank?
This is not from a book; it is from a handout.
Thank you for any help. I appreciate it.

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE AND TRY
Could you please help me
with this word problem. I'm really having a hard time
with it: Pipes A and B can fill a storage tank in 8
hours and 12 hours. With the tank empty, pipe A was
turned on at noon, and then pipe B was turned on at
1:30 P.M. At what time was the tank full?
Thank you.
1 solutions
Answer 13931 by venugopalramana(1619)on
2006-01-30 05:52:42 
Pipes A and B can fill a storage tank in 8 hours and
12 hours.
THAT IS PIPE A CAN FILL IN 1 HR.1/8 TANK AND
PIPE B CAN FILL IN 1 HOUR 1/12 TANK
With the tank empty, pipe A was turned on at noon, and
then pipe B was turned on at 1:30 P.M.
HENCE IN 1=30 PM - 12 = 00 NOON =1.5 HRS PIPE A FILLS
1.5*1/8 TANK=1.5/8 TANK
REST TO BE FILLED =1-1.5/8=6.5/8 TANK
WHEN A AND B ARE OPEN IN 1 HOUR THEY FILL 1/8 +1/12
TANK = 5/24 TANK
HENCE TO FILL 6.5/8 TANK THEY NEED
{(6.5/8)/(5/24)}=(24*6.5)/(8*5)=3.9 HRS.=3 HRS 54 MTS
HENCE THE TANK WILL BE FULL AT
1=30.....+.....3=54.......5=24 PM
At what time was the tank full?
The inlet pipe of an oil tank can fill the
tank in 1 hour, 30 minutes. The outlet pipe can empty
the tank in 1 hour. How long does it take to empty the
tank if both pipes are open?inlet pipe fills 1 tank in
1hr 30 mts =1.5=3/2 hrs.or inlet pipe fills
...1/(3/2)=2/3 tank in 1 hr.out let pipe empties 1
tank in 1 hr....so assuming the tank is full at the
start water to be emptied =1 tank and in 1 hr inlet
fills 2/3 tank and out let empties 1 tank ..so net out
flow is 1-2/3=(3-2)/3=1/3 tank in 1 hour.so 1/3 tank
emptying needs .....1 hr then 1 tank emptying needs
......? hrs...= 1*1/(1/3)=3 hrs.